OpenStudy (anonymous):

Find h if A = 80π and r = 5

OpenStudy (saifoo.khan):

is the question correct?

OpenStudy (anonymous):

* The formula for the surface area of a cylinder is A = 2πr(r + h). Find h if A = 80π and r = 5

OpenStudy (anonymous):

so sorry 3

OpenStudy (anonymous):

if u know the formula then u can solve this one and i nm noticing u test people

OpenStudy (saifoo.khan):

\[A = 2 \pi r ( r+ h)\] \[80 \pi = 2 \pi (5)(5+h)\] \[8 = 5 + h\] h = 8 - 5 h = 3

OpenStudy (anonymous):

Can you do this problem step by step with me to explain it?

OpenStudy (saifoo.khan):

\[A = 2 \pi r ( r+ h)\] \[80 \pi = 2 \pi (5)(5+h)\] \[80 \pi = 10 \pi (5+h)\] Divide 10 pi on both sides. \[\frac{80 \pi}{10 \pi} = (5+h)\] \[8 = 5 + h\] h = 8 - 5 h = 3